# The Case for Persistent CXL switches

### v1

@@ I Introduction
The emergence of Compute Express Link (CXL) interconnect technology is a significant event that addresses the ever increasing demand higher memory capacity in data center servers. With CXL, memory devices can be attached to the PCIe ports, accessible by the CPU via loads/stores instructions [1], allowing memory expansion over traditional (local) memory. Due to its packet-based protocol, CXL enables various types of memories to be attached, including volatile (DRAM) and byte-addressable non-volatile memory (Optane Pmem, memory-semantic SSD, etc.). From beginning, the CXL protocol was designed to be compatible with persistent memory (PM) devices, with features to manage persistency (e.g. global persistent flush) [2]. This is important because PM devices such as memory-semantic SSDs are projected to be much denser, hence provide a much larger capacity, than traditional DRAMs. Furthermore, many applications including databases and key-value stores, can benefit from the persistency in addition to the capacity [3].
@@
  • We evaluated PCS using Splash-4 parallel benchmark suite. Our evaluation shows that PCS significantly reduces the latency fo persist operation (up to 55%) resulting in 12% speedup, on average, of the evaluated workloads.
@@ IV-A Correctness Criteria
To ensure crash consistency, once data persistency is confirmed by any persistent structure, data must not leave persistent domain ( always a persistent copy of the latest version of the data exit in the persistent domain). if crash happen while data is out of persistent domain, data will be lost or get corrupted upon reboot.PCS must ensure there will be always a persistent copy of the the data in the persistent domain before sends acknowledgment of data persistency.
@@ VII Evaluation
Figure 5 shows PB and PB_RF achieve 12% and 15% average speedup over NoPB, respectively, for eight-threaded workload. The speedup is significant considering design simplicity and low hardware cost. For Radiosity and Lu_non, PB provides more than 20% speedup where PB_RF provides 40% speedup. However cholesky suffers 3% slowdown for PB and 13% slowdown for PB_RF.
@@
To assess the impact of PB and PB_RF on different workloads, we examine persist and read operation latencies. As shown in Figure 6(a) PB reduces persist latency by 43% to 56%, achieving the goal of PCS outlined in Section I. However, in PB_RF, persist latency increases for most workloads, with varying rates. For Radiocity, latency remains the same, while for Cholesky, it rises from 51% to 66% of NoPB. A similar increase in persist latency is observed for FFT. While PCS reduces persist latency as expected, its effect on overall workload performance is not straightforward (Figure 5).
Figure 6(b) explains why persist latency reduction doesn’t always lead to proportional performance improvement. Introducing PB increases read latency for all cases, as some reads are rerouted through the PI buffer (Section V-D3). Although the increase in read latency(2.5%-12%) is modest, workloads heavily reliant on read operations, like Cholesky (which has a 2.5% slowdown despite a 49% reduction in persist latency), experience overall performance degradation. FFT, however, does not suffer a performance loss (a 3% speedup despite a 50% increase in read latency). PB_RF reduces read latency compared to PB for most workloads, except Cholesky and Volrend_npl. For Cholesky, PB_RF increases read latency by 16% over NoPB, causing a 13% overall slowdown. Although FFT sees a reduction in read latency (from 50% to 44%), it still experiences a 2% slowdown due to increased persist latency.
@@
In Figure 7(a), we observe very high read hit rate at persist buffer for Radiocity (51%) compared to cholesky & Volrend_npl (around 1%). For other workload, this hit rate around 20%. For write coalescing (Figure 7(b)), we observe similar pattern except FFT which also have lower coalescing rate (2.8%). As Radiocity have highest read hit rate and write coalescing rate, it experience higher increase of speedup in PB_RF scheme over PB scheme. Where for cholesky & Volrend_npl very low read hit rate and write coalescing rate results in slowdown over PB scheme. As Lu_cont, Lu_non & Raytrace have noticeable read hit rate and write coalescing rate, they also demonstrate speedup grain for PB_RF over PB.
@@
Figure 8 shows speedup of Radiocity, cholesky & FFT over NoPB for different PBE count. Radiocity experienced high speedup for both PB and PB_RF, representing best case scenario while cholesky results in maximum slowdown for both representing worst case scenario. FFT experience speedup for PB but slowdown for PB_RF highlight contrasting effect between two scheme. With PBE count increase, tag access latency also increase which we calculated using cacti[25]. For PB, increasing PBE count does not show noticeable influence over speedup. For PB_RF scheme, both cholesky & FFT experience increasing slowdown with increase of PBE count, while cholesky experience at higher rate. Radiocity experience increasing speedup with increasing PBE count.

### v2

@@ I Introduction
The emergence of Compute Express Link (CXL) interconnect technology is a significant event that addresses the ever increasing demand for higher memory capacity in data center servers. With CXL, memory devices can be attached to the PCIe ports, accessible by the CPU via loads/stores instructions [1], allowing memory expansion over traditional (local) memory. Due to its packet-based protocol, CXL enables various types of memories to be attached, including volatile (DRAM) and byte-addressable non-volatile memory (Optane Pmem, memory-semantic SSD, etc.). From beginning, the CXL protocol was designed to be compatible with persistent memory (PM) devices, with features to manage persistency (e.g. global persistent flush) [2]. This is important because PM devices such as memory-semantic SSDs are projected to be much denser, hence provide a much larger capacity, than traditional DRAMs. Furthermore, many applications including databases and key-value stores, can benefit from the persistency in addition to the capacity [3].
@@
  • We evaluated PCS using Splash-4 parallel benchmark suite. Our evaluation shows that PCS significantly reduces the latency for persist operation (up to 55%) resulting in 12% speedup, on average, of the evaluated workloads.
@@ IV-A Correctness Criteria
To ensure crash consistency, once data persistency is confirmed by any persistent structure, data must not leave persistent domain ( always a persistent copy of the latest version of the data exit in the persistent domain). if crash happen while data is out of persistent domain, data will be lost or get corrupted upon reboot.PCS must ensure there will always be a persistent copy of the the data in the persistent domain before sends acknowledgment of data persistency.
@@ VII Evaluation
Figure 5 shows PB and PB_RF achieve 12% and 15% average speedup over NoPB, respectively, for eight-threaded workload. The speedup is significant considering design simplicity and low hardware cost. For Radiosity and Lu_non, PB provides more than 20% speedup where PB_RF provides 40% speedup. However Choelesky suffers 3% slowdown for PB and 13% slowdown for PB_RF.
@@
To assess the impact of PB and PB_RF on different workloads, we examine persist and read operation latencies. As shown in Figure 6(a) PB reduces persist latency by 43% to 56%, achieving the goal of PCS outlined in Section I. However, in PB_RF, persist latency increases for most workloads, with varying rates. For Radiocity, latency remains the same, while for Choelesky, it rises from 51% to 66% of NoPB. A similar increase in persist latency is observed for FFT. While PCS reduces persist latency as expected, its effect on overall workload performance is not straightforward (Figure 5).
Figure 6(b) explains why persist latency reduction doesn’t always lead to proportional performance improvement. Introducing PB increases read latency for all cases, as some reads are rerouted through the PI buffer (Section V-D3). Although the increase in read latency(2.5%-12%) is modest, workloads heavily reliant on read operations, like Choelesky (which has a 2.5% slowdown despite a 49% reduction in persist latency), experience overall performance degradation. FFT, however, does not suffer a performance loss (a 3% speedup despite a 50% increase in read latency). PB_RF reduces read latency compared to PB for most workloads, except Choelesky and Volrend_npl. For Choelesky, PB_RF increases read latency by 16% over NoPB, causing a 13% overall slowdown. Although FFT sees a reduction in read latency (from 50% to 44%), it still experiences a 2% slowdown due to increased persist latency.
@@
In Figure 7(a), we observe very high read hit rate at persist buffer for Radiocity (51%) compared to Choelesky & Volrend_npl (around 1%). For other workload, this hit rate around 20%. For write coalescing (Figure 7(b)), we observe similar pattern except FFT which also have lower coalescing rate (2.8%). As Radiocity have highest read hit rate and write coalescing rate, it experience higher increase of speedup in PB_RF scheme over PB scheme. Where for Choelesky & Volrend_npl very low read hit rate and write coalescing rate results in slowdown over PB scheme. As Lu_cont, Lu_non & Raytrace have noticeable read hit rate and write coalescing rate, they also demonstrate speedup gain for PB_RF over PB.
@@
Figure 8 shows speedup of Radiocity, Choelesky & FFT over NoPB for different PBE count. Radiocity experienced high speedup for both PB and PB_RF, representing best case scenario while Choelesky results in maximum slowdown for both representing worst case scenario. FFT experience speedup for PB but slowdown for PB_RF highlight contrasting effect between two scheme. With PBE count increase, tag access latency also increase which we calculated using cacti[25]. For PB, increasing PBE count does not show noticeable influence over speedup. For PB_RF scheme, both Choelesky & FFT experience increasing slowdown with increase of PBE count, while Choelesky experience at higher rate. Radiocity experience increasing speedup with increasing PBE count.